### v2

# Dynamics of Dehn Twists in the Outer Automorphism Group of a Free Group

Abstract.
We study Dehn twists in the outer automorphism group of a finitely generated non-abelian free group. Our main result states that, under certain compatibility conditions, sufficiently large powers of finitely many Dehn twists generate a right-angled Artin group. The proof proceeds by analyzing the geometry of spheres, tori, and simple closed curves in a doubled handlebody. Along the way, we establish the bigon–bihedron criterion and an equivalent condition for commuting Dehn twists. Furthermore, we construct a compact topological space on which every Dehn twist acts parabolically.
Key words and phrases:
outer automorphism group of free group, doubled handlebody, right-angled Artin group
1991 Mathematics Subject Classification:
primary: 57M07, 57M15, 20F28, secondary: 20F36, 57M60
1. Introduction
Let $\mathbb{F}=\mathbb{F}_{g}$ be a free group of rank $g$. An $\mathbb{F}$-tree is a simplicial tree on which $\mathbb{F}$ acts simplicially without edge inversions. An $\mathbb{F}$-tree is called a $\mathbb{Z}$-splitting if every edge stabilizer is infinite cyclic. It is said to be one-edge if the induced action of $\mathbb{F}$ on the set of edges is transitive. In this case, the conjugacy class of an edge stabilizer of a one-edge $\mathbb{Z}$-splitting $T$ is uniquely determined, and we call it the core of $T$.
For a conjugacy class $\gamma$ of an infinite cyclic subgroup of $\mathbb{F}$, the intersection number of $\gamma$ and $T$, denoted by $i(\gamma,T)$, is defined as the minimal translation length of a nontrivial element of a subgroup representing $\gamma$ on $T$:
$i(\gamma,T)=\min\{\,d_{T}(x,wx)\mid 1\neq w\in c\in\gamma,\;x\in T\,\},$
where $d_{T}$ denotes the combinatorial metric on $T$. For example, if $\gamma$ is the core of $T$, then $i(\gamma,T)=0$.
The automorphism group $\operatorname{Aut}(\mathbb{F})$ contains the inner automorphism group $\operatorname{Inn}(\mathbb{F})$ as a normal subgroup, and the quotient
$\operatorname{Out}(\mathbb{F}):=\operatorname{Aut}(\mathbb{F})/\operatorname{Inn}(\mathbb{F})$
is called the outer automorphism group of $\mathbb{F}$. Given a one-edge $\mathbb{Z}$-splitting $T$ of $\mathbb{F}$, one can associate a Dehn twist (or generalized transvection) $\delta_{\vec{T}}\in\operatorname{Out}(\mathbb{F})$, defined as an outer automorphism induced by twisting along $T$.
Our main theorem concerns the existence of right-angled Artin subgroups generated by powers of such Dehn twists. For a finite simplicial graph $\Gamma=(V,E)$, the right-angled Artin group (RAAG) on $\Gamma$ is
$A(\Gamma):=\langle\,v\in V\mid[u,v]=1\text{ whenever }\{u,v\}\in E\,\rangle.$
Given a finite collection $\{T_{1},\dots,T_{k}\}$ of one-edge $\mathbb{Z}$-splittings of $\mathbb{F}$, one defines the coincidence graph $\Gamma$ with vertices $T_{1},\dots,T_{k}$, where $T_{i}$ and $T_{j}$ are adjacent precisely when they admit a common refinement. For integers $n_{1},\dots,n_{k}$, this determines a homomorphism
$A(\Gamma)\longrightarrow\operatorname{Out}(\mathbb{F}),\qquad T_{i}\longmapsto\delta_{\vec{T}_{i}}^{\,n_{i}}.$
Our main result asserts that, under suitable hypotheses, this map is injective and its image is a right-angled Artin subgroup of $\operatorname{Out}(\mathbb{F})$.
Theorem 1.1 (Theorem 5.5).
Let $V=\{T_{1},\dots,T_{k}\}$ be a compatible collection of one-edge $\mathbb{Z}$-splittings with cores $\gamma_{1},\dots,\gamma_{k}$, and let $\Gamma$ be the coincidence graph of $V$. For each $i$, let $\delta_{i}=\delta_{\vec{T}_{i}}$ be a Dehn twists along $T_{i}$. Then for all
$\lvert n_{1}\rvert,\dots,\lvert n_{k}\rvert\geq 5\,\max_{i\neq j}{i(\gamma_{i},T_{j})}+8,$
the homomorphism
$\Phi:A(\Gamma)\longrightarrow\operatorname{Out}(\mathbb{F}),\qquad T_{i}\mapsto\delta_{i}^{n_{i}}$
is injective.
The problem of constructing free subgroups of $\operatorname{Out}(\mathbb{F})$ has been extensively studied. Bestvina, Feighn, and Handel [1] proved that $\operatorname{Out}(\mathbb{F})$ satisfies the Tits alternative: every subgroup is either virtually solvable or contains a free subgroup of rank two. Kapovich and Lustig [13] showed that sufficiently large powers of two hyperbolic iwip elements generate either a virtually cyclic subgroup or a nonabelian free group. Uyanik [19] obtained a related result in the setting of hyperbolic extensions. Clay and Pettet [4] proved that if two one-edge $\mathbb{Z}$-splittings are filling, then sufficiently large powers of the associated Dehn twists generate a free group of rank two consisting of atoroidal fully irreducible elements except for the conjugacy classes of the twists themselves. Gültepe [9] provided another criterion under which powers of two Dehn twists generate a free subgroup of rank two.
A complementary development is due to Taylor [18], who proved that every right-angled Artin group embeds quasi-isometrically into some $\operatorname{Out}(\mathbb{F})$. He constructed a suitable collection of fully irreducible outer automorphisms supported on free factors, whose sufficiently large powers generate the desired RAAG. Since these automorphisms are fully irreducible, Dehn twists do not appear in his construction. Our work may therefore be viewed as a counterpart to Taylor’s theorem: we construct right-angled Artin subgroups of $\operatorname{Out}(\mathbb{F})$ generated instead by powers of Dehn twists.
Theorem 1.1 gives a broad and explicit method for constructing RAAG subgroups inside $\operatorname{Out}(\mathbb{F})$. However, the compatibility condition does not encompass all right-angled Artin phenomena arising from Dehn twists. For instance, $\operatorname{Out}(\mathbb{F}_{g})$ contains a direct product of 2g42g-4 free groups generated by Dehn twists (Horbez and Wade [11], and Bridson and Wade [3]) and a free-abelian subgroup of rank 2g32g-3 generated by Dehn twists (Feighn and Handel [7]), yet the associated one-edge $\mathbb{Z}$–splittings in these examples are not compatible. Thus, Theorem 1.1 captures a large and well-structured class of RAAGs generated by powers of Dehn twists, but it does not apply to every such instance found in $\operatorname{Out}(\mathbb{F}_{g})$.
We approach the proof of Theorem 1.1 by working in the mapping class group of the doubled handlebody
$\mathbb{M}=\mathbb{M}_{g}=\#_{g}(\mathbb{S}^{1}\times\mathbb{S}^{2}).$
The interaction between $\operatorname{Out}(\mathbb{F})$ and $\operatorname{MCG}(\mathbb{M})$ was first studied by Whitehead [20, 21] in his solution to the automorphic equivalence problem for $\mathbb{F}$. In particular, he proved that the natural map
$\operatorname{MCG}(\mathbb{M})\longrightarrow\operatorname{Out}(\mathbb{F})$
is surjective. Laudenbach [15] later established the short exact sequence
$1\longrightarrow(\mathbb{Z}/2\mathbb{Z})^{g}\longrightarrow\operatorname{MCG}(\mathbb{M})\longrightarrow\operatorname{Out}(\mathbb{F})\longrightarrow 1.$
Brendle, Broaddus, and Putman [2] subsequently proved that this sequence splits; in particular, there exists a section
$\operatorname{Out}(\mathbb{F})\longrightarrow\operatorname{MCG}(\mathbb{M}).$
In this setting, a Dehn twist may be interpreted as the isotopy class of a homeomorphism obtained by twisting $\mathbb{M}$ along an essential torus. Gültepe [9] likewise studied Dehn twists realized as twists along essential tori.
We begin by analyzing pairs of commuting Dehn twists. Given homotopy classes of essential tori $T$ and $T^{\prime}$, their geometric intersection number $i(T,T^{\prime})$ is defined as the minimum number of intersection circles among all representatives $t\in T$ and $t^{\prime}\in T^{\prime}$. If $i(T,T^{\prime})=0$, that is, if $T$ and $T^{\prime}$ admit disjoint representatives, then the corresponding Dehn twists commute. Our next result establishes the converse: two Dehn twists commute if and only if their supports may be chosen disjointly. In particular, no accidental commutation occurs.
Theorem 1.2 (Theorem 3.7).
Let $T$ and $T^{\prime}$ be homotopy classes of essential tori. Then the following statement holds:
$\delta_{\vec{T}},\delta_{\vec{T}^{\prime}}]=1$ if and only if $i(T,T^{\prime})=0$.
From an algebraic perspective, two Dehn twists commute precisely when their associated one-edge $\mathbb{Z}$-splittings admit a common refinement. On the geometric side, recall that two simple closed curves on a surface intersect minimally if and only if they form no bigon—the classical bigon criterion, which underlies much of surface topology. In an analogous spirit, our work is grounded in the following criterion governing intersections of essential tori in doubled handlebodies.
Proposition 1.3 (Bigon–bihedron criterion, Proposition 3.4).
Two essential tori intersect minimally if and only if they form neither a bigon nor a bihedron.
One can formulate analogous criteria in the cases of two essential spheres or of an essential sphere together with an essential torus. In addition, one obtains a bigon criterion in this setting: an essential simple closed curve and an essential sphere (or an essential torus) intersect minimally if and only if they do not form a bigon. We note that Hatcher [10] introduced a normal form for sphere systems, and this normal form guarantees that any two sphere systems placed in it are in minimal position.
Meanwhile, Theorem 1.1 is proved by a ping–pong argument on the set of homotopy classes of essential simple closed curves. We employ a variation of the ping–pong method due to Koberda [14]. Careful control of intersection numbers within this framework yields the proof of Theorem 1.1.
In addition, we construct a compact topological space $\overline{(p\circ i_{*})(\mathcal{C})}$ on which $\operatorname{Out}(\mathbb{F})$ acts continuously. This space is precisely the closure of the set of points represented by homotopy classes of essential simple closed curves that represent $1$-free factors of $\mathbb{F}$. The definition of $\overline{(p\circ i_{*})(\mathcal{C})}$ is closely analogous to that of the projective space $\mathcal{PML}(S)$ of measured laminations on a surface $S$. In $\mathcal{PML}(S)$, the action of a Dehn twist is parabolic, and a Dehn twist in $\operatorname{Out}(\mathbb{F})$ exhibits the same dynamical behavior on $\overline{(p\circ i_{*})(\mathcal{C})}$, as described below.
Theorem 1.4 (Theorem 6.5).
Let $T$ be a homotopy class of an essential torus with core $\gamma$. Let $\alpha$ be a homotopy class of an essential simple closed curve that represents a free factor. Then, we have $i(\alpha,T)>0$ if and only if $\delta_{\vec{T}}^{n}(\alpha)$ converges to $\gamma$ as $n\to\infty$.
1.1. Acknowledgments
The author thanks Sang-hyun Kim for many helpful and inspiring conversations during the preparation of this work. During the revision process, discussions with Michael Kopreski, Thomas Hill, George Shaji, and Brian Udall highlighted an error in a previous version of Lemma 3.6, leading to its improvement. The author is also grateful to Thomas Koberda for valuable comments that strengthened the manuscript. Thanks to Ignat Soroko, the author became aware of Whitehead’s work.
This paper is dedicated to my son, Jian, who arrived as this manuscript was taking shape.
2. Dehn twist
In this section, we review the basic background on outer automorphism groups of free groups and doubled handlebodies required for our discussion. Let $\mathbb{F}=\mathbb{F}_{g}$ be a free group of rank $g\geq 2$. We write $\operatorname{Aut}(\mathbb{F})$ and $\operatorname{Inn}(\mathbb{F})$ for the automorphism group and the inner automorphism group of $\mathbb{F}$, respectively. Note $\operatorname{Inn}(\mathbb{F})$ is a normal subgroup of $\operatorname{Aut}(\mathbb{F})$. The outer automorphism group of $\mathbb{F}$, denoted by $\operatorname{Out}(\mathbb{F})$, is the quotient of $\operatorname{Aut}(\mathbb{F})$ by $\operatorname{Inn}(\mathbb{F})$.
2.1. Dehn twist automorphisms of free groups
An $\mathbb{F}$-tree $T$ is called a one-edge $\mathbb{Z}$-splitting if the action of $\mathbb{F}$ on the set of edges is transitive and a stabilizer of an edge is an infinite cyclic subgroup of $\mathbb{F}$. Let $\vec{E}T$ be the set of orientations of edges of $T$. Then there exists an orientation reversing $\vec{E}T\to\vec{E}T$, that is, $(\vec{e})$ is the reverse orientation of $\vec{e}$ for each $\vec{e}\in\vec{E}T$. Let $o:\vec{E}T\to\mathbb{F}$ be a map defined by
  1. (1)
    $\langle o(\vec{e})\rangle=\operatorname{Stab}(e)$,
  2. (2)
    $o(*(\vec{e}))=o(\vec{e})^{-1}$ and
  3. (3)
    $o(w\vec{e})=wo(\vec{e})w^{-1}$
for all $\vec{e}\in\vec{E}T$ and $w\in\mathbb{F}$. We call $o$ an orientation of $T$ and write $\vec{T}:=(T,o)$ for an oriented one-edge $\mathbb{Z}$-splitting.
Fix a base-vertex $v_{0}\in T$. For each vertex $v\in T$, there exists a unique geodesic path from $v_{0}$ to $v$, denoted by $\vec{e}_{1},\dots,\vec{e}_{k}$. Then we define
$\delta_{\vec{T},v_{0}}(v):=o(\vec{e}_{1})\dots o(\vec{e}_{k})v,$
which is called a Dehn twist automorphism based at $v_{0}$. We then show the following.
Lemma 2.1.
Let $\delta_{\vec{T},v_{0}}:VT\to VT$ be defined as above. Then $\delta_{\vec{T},v_{0}}$ can be extended to a simplicial automorphism of $T$.
Proof.
Suppose $v$ and $v^{\prime}$ are adjacent. Without loss of generality, let us assume that $v^{\prime}$ is closer to $v_{0}$ than $v$. Let $\vec{e}_{1},\dots,\vec{e}_{k}$ be the geodesic path from $v_{0}$ to $v$. Since $o(\vec{e}_{k})$ stabilizes the edge joining $v^{\prime}$ and $v$, we have
$d_{T}(\delta_{\vec{T},v_{0}}(v),\delta_{\vec{T},v_{0}}(v^{\prime}))$ $=d_{T}(o(\vec{e}_{1})\dots o(\vec{e}_{k-1})o(\vec{e}_{k})v,o(\vec{e}_{1})\dots o(\vec{e}_{k-1})v^{\prime})$
$=d_{T}(o(\vec{e}_{k})v,v^{\prime})=1.$
That is, $\delta_{\vec{T},v_{0}}(v)$ and $\delta_{\vec{T},v_{0}}(v^{\prime})$ are adjacent. So $\delta_{\vec{T},v_{0}}$ can be extended to a simplicial automorphism of $T$. ∎
Lemma 2.2.
For each vertex $v\in T$, if $\vec{e}_{1},\dots,\vec{e}_{k}$ is the geodesic path from $v_{0}$ to $v$, let us define
$\delta_{\vec{T},v_{0}}^{-1}(v):=o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k})^{-1}(v).$
Then the map $\delta_{\vec{T},v_{0}}^{-1}:VT\to VT$ is the inverse of $\delta_{\vec{T},v_{0}}$.
Proof.
Since the sequence
$\vec{e}_{1},o(\vec{e}_{1})^{-1}\vec{e}_{2},\dots,o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k-1})^{-1}\vec{e}_{k}$
is the geodesic path from $v_{0}$ to $o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k})^{-1}v$, we obtain
$\delta_{\vec{T},v_{0}}\delta_{\vec{T},v_{0}}^{-1}(v)$ $=\delta_{\vec{T},v_{0}}(o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k})^{-1}v)$
$=o(\vec{e}_{1})o(o(\vec{e}_{1})^{-1}\vec{e}_{2})\dots o(o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k-1})^{-1}\vec{e}_{k})(o(\vec{e}_{1})^{-1}\dots o(\vec{e}_{k})^{-1}v)$
$=v.$
Note $\vec{e}_{1},o(\vec{e}_{1})\vec{e}_{2},\dots,o(\vec{e}_{1})\dots o(\vec{e}_{k-1})\vec{e}_{k}$ is the geodesic path from $v_{0}$ to $\delta_{\vec{T},v_{0}}(v)$. So we have
$\delta_{\vec{T},v_{0}}^{-1}\delta_{\vec{T},v_{0}}(v)$ $=\delta_{\vec{T},v_{0}}^{-1}(o(\vec{e}_{1})\dots o(\vec{e}_{k})v)$
$=o(\vec{e}_{1})^{-1}o(o(\vec{e}_{1})\vec{e}_{2})^{-1}\dots o(o(\vec{e}_{1})\dots o(\vec{e}_{k-1})\vec{e}_{k})^{-1}(o(\vec{e}_{1})\dots o(\vec{e}_{k})v)$
$=v$
Therefore, the statement holds. ∎
Lemma 2.3.
For each $w\in\mathbb{F}$, it holds that $w\delta_{\vec{T},v_{0}}w^{-1}=\delta_{\vec{T},wv_{0}}$.
Proof.
For a vertex $v\in VT$, let $\vec{e}_{1},\dots,\vec{e}_{k}$ be the geodesic from $wv_{0}$ to $v$. Since $w^{-1}\vec{e}_{1},\dots,w^{-1}\vec{e}_{k}$ is a geodesic from $v_{0}$ to $w^{-1}v$, we have
$w\delta_{\vec{T},v_{0}}w^{-1}v=wo(w^{-1}\vec{e}_{1})\dots o(w^{-1}\vec{e}_{k})w^{-1}v=o(\vec{e}_{1})\dots o(\vec{e}_{k})v=\delta_{\vec{T},wv_{0}}(v)$
So the statement holds. ∎
Lemma 2.4.
For a vertex $v\in T$, if $\vec{e}_{1},\dots,\vec{e}_{k}$ is any path from $v_{0}$ to $v$, then we have $\delta_{\vec{T},v_{0}}(v)=o(\vec{e}_{1})\dots o(\vec{e}_{k})v$.
Proof.
If $\vec{e}_{1},\dots,\vec{e}_{k}$ is not geodesic, then there exists $i$ such that $\vec{e}_{i+1}=*(\vec{e}_{i})$. So we have $o(\vec{e}_{1})\dots o(\vec{e}_{k})=o(\vec{e}_{1})\dots o(\vec{e}_{i-1})o(\vec{e}_{i+2})\dots o(\vec{e}_{k})$, which corresponds to the path $\vec{e}_{1},\dots,\vec{e}_{i-1},\vec{e}_{i+2},\dots,\vec{e}_{k}$. So we can inductively reduce the length of the path until it is a geodesic. Therefore, we get $o(\vec{e}_{1})\dots o(\vec{e}_{k})v=\delta_{\vec{T},v_{0}}(v)$. ∎
Proposition 2.5.
For each $w\in\mathbb{F}$, we have $\delta_{\vec{T},v_{0}}w\delta_{\vec{T},v_{0}}^{-1}\in\mathbb{F}$.
Proof.
Let $\vec{e}_{1},\dots,\vec{e}_{k}$ be the geodesic path from $v_{0}$ to $wv_{0}$. For a vertex $v\in T$, if $\vec{e}_{1}^{\prime},\dots,\vec{e}_{m}^{\prime}$ is a geodesic path from $v_{0}$ to $v$, then we have
$\delta_{\vec{T},v_{0}}w\delta_{\vec{T},v_{0}}^{-1}(v)$ $=\delta_{\vec{T},v_{0}}(w\delta_{\vec{T},v_{0}}^{-1}w^{-1})(wv)$
$=\delta_{\vec{T},v_{0}}\delta_{\vec{T},wv_{0}}^{-1}(wv)$
$=\delta_{\vec{T},v_{0}}(o(w\vec{e}_{1}^{\prime})^{-1}\dots o(w\vec{e}_{m}^{\prime})^{-1}(wv))$
$=o(\vec{e}_{1})\dots o(\vec{e}_{k})\cdot$
$\quad\,\,o(w\vec{e}_{1}^{\prime})o(o(w\vec{e}_{1}^{\prime})^{-1}w\vec{e}_{2}^{\prime})\dots o(o(w\vec{e}_{1}^{\prime})^{-1}\dots o(w\vec{e}_{m-1}^{\prime})^{-1}w\vec{e}_{m})\cdot$
$\quad\,\,o(w\vec{e}_{1}^{\prime})^{-1}\dots o(w\vec{e}_{m}^{\prime})^{-1}(wv)$
$=o(\vec{e}_{1})\dots o(\vec{e}_{k})wv.$
Therefore, we conclude that $\delta_{\vec{T},v_{0}}w\delta_{\vec{T},v_{0}}^{-1}=o(\vec{e}_{1})\dots o(\vec{e}_{k})w\in\mathbb{F}$. ∎
So one considers the conjugation by $\delta_{\vec{T},v_{0}}$ on $\mathbb{F}$ as an automorphism of $\mathbb{F}$. Then the outer automorphism represented by the conjugation by $\delta_{\vec{T},v_{0}}$ is called a Dehn twist, denoted by $\delta_{\vec{T}}$.
2.2. Dehn twists of doubled handlebodies
Let $\mathbb{M}=\mathbb{M}_{g}$ denote the connected sum of $g$ copies of $\mathbb{S}^{1}\times\mathbb{S}^{2}$, namely
$\mathbb{M}=\mathbb{M}_{g}=\#_{g}\,\mathbb{S}^{1}\times\mathbb{S}^{2}=\underbrace{(\mathbb{S}^{1}\times\mathbb{S}^{2})\#\dots\#(\mathbb{S}^{1}\times\mathbb{S}^{2})}_{g~\text{times}}.$
The fundamental group $\pi_{1}(\mathbb{M},x)$, for a chosen basepoint $x\in\mathbb{M}$, is a free group of rank $g$. We henceforth identify $\mathbb{F}$ with $\pi_{1}(\mathbb{M},x)$.
Let $\operatorname{MCG}(\mathbb{M})$ denote the mapping class group of $\mathbb{M}$, that is, the group of isotopy classes of orientation-preserving homeomorphisms. Laudenbach [15] shows the following short exact sequence holds:
$1\longrightarrow(\mathbb{Z}/2\mathbb{Z})^{g}\longrightarrow\operatorname{MCG}(\mathbb{M})\longrightarrow\operatorname{Out}(\mathbb{F})\longrightarrow 1$
Brendle, Broaddus and Putman [2] discover that this sequence is splitting, that is, there exists a section $\operatorname{Out}(\mathbb{F})\hookrightarrow\operatorname{MCG}(\mathbb{M})$. It is therefore natural to refer to the image of a Dehn twist in $\operatorname{MCG}(\mathbb{M})$ simply as a Dehn twist. What kind of a homeomorphism is associated with a Dehn twist? To answer to this question, we need to dive into the geometry of $\mathbb{M}$.
For a torus $t$ embedded in $\mathbb{M}$, we say $t$ is essential if
  1. (1)
    it does not bound a solid torus and
  2. (2)
    there exists a simple closed curve in $t$ which is essential in $\mathbb{M}$.
By the definition, an induced map from a fundamental group of an essential torus $t$ to $\mathbb{F}$ has an infinite cyclic image. Thus, the simple closed curve in (2) is uniquely determined up to homotopy. We refer to the homotopy class of this curve as the core of $t$.
Let $\widetilde{\mathbb{M}}$ be a universal cover of $\mathbb{M}$ with a covering map $\pi:\widetilde{\mathbb{M}}\to\mathbb{M}$. For an essential torus $t\subset\mathbb{M}$, the dual tree of $t$ is a tree $T$ embedded in $\widetilde{\mathbb{M}}$ satisfying
  1. (1)
    each component of $\widetilde{\mathbb{M}}\setminus\pi^{-1}(t)$ contains exactly one vertex of $T$ and
  2. (2)
    each component of $\pi^{-1}(t)$ intersects exactly one edge of $T$.
For each $v\in VT$ and $w\in\mathbb{F}$, if $C$ is the component of $\widetilde{\mathbb{M}}$ that contains $v$, we define $wv$ as the vertex contained in $wC$. This gives a simplicial action of $\mathbb{F}$ on $T$, which makes $T$ a one-edge $\mathbb{Z}$-splitting.
For an orientation on $T$, the Dehn twist $\delta_{\vec{T}}$ along $\vec{T}$ is associated with some mapping class of $\operatorname{MCG}(\mathbb{M})$ through the section $\operatorname{Out}(\mathbb{F})\hookrightarrow\operatorname{MCG}(\mathbb{M})$. This is represented by the following homeomorphism.
Let $t$ be an essential torus, and let $N(t)$ denote a closed regular neighborhood of $t$. Choose an orientation on the core of $t$, and denote by $\vec{t}$ the torus $t$ equipped with this orientation.
Note $t$ is homeomorphic to $\mathbb{S}^{1}\times\mathbb{S}^{1}$. Assume $\mathbb{S}^{1}$ is oriented. There exists a homeomorphism
$h:N(t)\to[0,1]\times\mathbb{S}^{1}\times\mathbb{S}^{1}$
such that the restriction
$h^{-1}\big|_{\{1/2\}\times\{0\}\times\mathbb{S}^{1}}$
is orientation-preserving between $\{1/2\}\times\{0\}\times\mathbb{S}^{1}$ and the core of $t$. Define a map
$\delta:[0,1]\times\mathbb{S}^{1}\times\mathbb{S}^{1}\to[0,1]\times\mathbb{S}^{1}\times\mathbb{S}^{1}$
by
$\delta(t,s_{0},s_{1})=(t,s_{0},s_{1}+t)$
for all $(t,s_{0},s_{1})\in[0,1]\times\mathbb{S}^{1}\times\mathbb{S}^{1}$. Then the Dehn twist homeomorphism along $\vec{t}$, denoted $\delta_{\vec{t}}$, is defined by
$\delta_{\vec{t}}(x)\;=\;\begin{cases}h^{-1}\,\delta\,h(x),&x\in N(t),\\ x,&\text{otherwise.}\end{cases}$
Lemma 2.6 ([9, Lemma 2.6]).
Let $t$ be an essential torus of $\mathbb{M}$. Let $T$ be the dual tree of $t$. For each orientation of $t$, there exists an orientation of $T$ such that $\delta_{\vec{t}}$ represents $\delta_{\vec{T}}$.
Note if two essential tori are homotopic, then they give the same Dehn twist as outer automorphisms. By Brendle, Broaddus and Putman [2], this outer automorphism can be considered as a mapping class. Note the mapping class group consists of isotopy classes of orientation-preserving homeomorphisms. So the following holds.
Proposition 2.7.
Let $\vec{t}$ and $\vec{t}^{\prime}$ be oriented essential tori. Then, $\delta_{\vec{t}}$ is isotopic to $\delta_{\vec{t}^{\prime}}$ if and only if $\vec{t}$ is homotopic to $\vec{t}^{\prime}$.
We postpone the proof of Proposition 2.7 until after Lemma 3.6. Throughout this section, we will abuse notation by writing $T$ both for a homotopy class of an essential torus $t$ and for the associated dual tree of $t$. Similarly, we use $\vec{T}$ to denote either oriented version. With this convention, the Dehn twist along $\vec{T}$ will be denoted by $\delta_{\vec{T}}$.
3. Intersection number
Let $A$ and $B$ be homotopy classes of embedded submanifolds (e.g. simple closed curves, spheres and tori) of $\mathbb{M}$. The intersection number of $A$ and $B$, denoted by $i(A,B)$, is defined by
$i(A,B):=\min\{\lvert\pi_{0}(a\cap b)\rvert\mid a\in A,b\in B\}$
where $\pi_{0}(a\cap b)$ denotes the collection of components of $a\cap b$, that is, the zeroth homotopy group of $a\cap b$.
Example 3.1.
If $A$ and $B$ are homotopy classes of essential simple closed curves, then $i(A,B)$ is always zero. If $A$ is a homotopy class of an essential simple closed curve and $B$ is a homotopy class of an essential torus or an essential sphere, then $i(A,B)$ can be written by $\min\{\lvert a\cap b\rvert\mid a\in A,b\in B\}$. If $A$ and $B$ are homotopy classes of essential spheres and essential tori, then for each $a\in A$ and $b\in B$, the intersection of $a$ and $b$ is a disjoin union of circles, so $i(A,B)$ indicates the minimum number of circles among such intersections.
3.1. Bigon, bihedron and minimal position
Let $c$ be an essential simple closed curve, and let $s$ denote an essential sphere or an essential torus. We say $c$ and $s$ form a bigon if there are arcs $a\subset c$ and $a^{\prime}\subset s$ such that $a\cup a^{\prime}$ is a closed curve that bounds a disk. If $c$ and $s$ form a bigon, one can reduce the number of bigons by homotoping $c$. This induces the cardinality of $c\cap s$ is strictly larger than the intersection number of $[c]$ and $[s]$. What we want to mention is that the converse is also true.
Proposition 3.2 (Bigon criterion).
Let $c$ be an essential simple closed curve, and let $s$ be an essential sphere or an essential torus. Then, $\lvert c\cap s\rvert=i([c],[s])$ if and only if $c$ and $s$ do not form a bigon.
The following proof is adapted from Farb and Margalit [5].
Proof.
Suppose $\lvert c\cap s\rvert>i([c],[s])$. Let $c^{\prime}$ be a simple closed curve that is homotopic to $c$ such that $\lvert c^{\prime}\cap s\rvert=i([c],[s])$. Then there exists a homotopy $h$ that moves $c$ to $c^{\prime}$. Note the domain of $h$ is an annulus $A$. Let $\partial_{0}$ be the boundary component of $A$ whose image is $c$.
Note $h^{-1}(s)$ is a union of arcs whose endpoints are contained in the boundary of $A$. Because $c\cap s$ is strictly larger than $c^{\prime}\cap s$, there exists an arc $a$ of $h^{-1}(s)$ that connects two points of $\partial_{0}$. Then one can find a subarc $a_{0}\subset\partial_{0}$ such that $a\cup a_{0}$ bounds a disk in $A$. That is, the union of $h(a)$ and $h(a_{0})$ bounds a disk. Hence, $c$ and $s$ form a bigon. ∎
We may focus on an intersection number of two essential tori. There are two ways to reduce the number of circles of two essential tori via a homotopy.
Example 3.3.
Let $t$ and $t^{\prime}$ be essential tori.
  1. (1)
    Suppose there exist disks $D\subset t$ and $D^{\prime}\subset t^{\prime}$ such that $D\cup D^{\prime}$ forms a sphere bounding a $3$-ball. In this case, $t$ and $t^{\prime}$ are said to form a bihedron. By homotoping $t$ across this $3$-ball, one can eliminate the intersection circle $D\cap D^{\prime}$, thereby reducing the number of intersection circles between $t$ and $t^{\prime}$.
  2. (2)
    Suppose there are arcs $a\subset t$ and $a^{\prime}\subset t^{\prime}$ connecting two distinct intersection circles such that $a\cup a^{\prime}$ bounds a disk. Then $t$ and $t^{\prime}$ are said to form a bigon. By homotoping $t$ across this disk, one can reduce the number of intersection circles.
The above example applies both to pairs of essential spheres and to pairs consisting of an essential sphere and an essential torus. This naturally raises the question of whether forming a bigon or a bihedron is the only obstruction to realizing the intersection number. The following result gives a positive answer, which we refer to as the bigon–bihedron criterion.
Proposition 3.4 (Bigon–bihedron criterion).
Let $s$ and $t$ be essential spheres, essential tori, or one of each. Then, the number of intersection circles of $s$ and $t$ is exactly $i([s],[t])$ if and only if $s$ and $t$ form neither a bigon nor a bihedron.
Proof.
If $s$ and $t$ form either a bigon or a bihedron, then the number of intersection circles can be reduced, as in the preceding example.
Conversely, suppose $s$ and $t$ are not in minimal position. Then there exists a homotopy
$H:\hat{s}\times[0,1]\to\mathbb{M}$
[…]
The preimage $H^{-1}(t)$ is a (possibly disconnected) orientable surface with boundary. If $H^{-1}(t)$ contains a disk whose boundary lies in $s$, then $s\cup t$ contains a bihedron. Otherwise, since the number of circles in $s^{\prime}\cap t$ is strictly smaller than in $s\cap t$, there exists a connected component $S\subset H^{-1}(t)$ such that $S\cap(\hat{s}\times\{0\})$ is a disjoint union of at least two circles.
Choose two such circles, denoted by $C_{1}$ and $C_{2}$. Then there exist arcs $a\subset\hat{s}\times{0}$ and $b\subset S$, each connecting $C_{1}$ to $C_{2}$, such that the union $a\cup b$ bounds a disk in $\hat{s}\times[0,1]$. Since the union $H(a)\cup H(b)$ also bounds a disk, it follows that $s$ and $t$ form a bigon.
Therefore, if two essential tori are not in minimal position, they necessarily form either a bigon or a bihedron. ∎
By Laudenbach [15], two homotopic sphere systems are isotopic. Hatcher [10] show that two sphere systems can be in “normal form.” This is equivalent to minimal position. That is, two sphere systems are in normal form if and only if they are in minimal position. Refer to Horbez’s article [12].
The situation is not different if we consider an essential torus and a sphere system. Gültepe [8] defines the normal form of a torus with respect to a sphere system. This is equivalent to minimal position.
Theorem 3.5 (Minimal position).
Let $A_{1},\dots,A_{k}$ be homotopy classes of essential simple closed curves, essential spheres and essential tori. Then there are representatives
$a_{1}\in A_{1},\dots,a_{k}\in A_{k}$
such that
$\lvert\pi_{0}(a_{i}\cap a_{j})\rvert=i(A_{i},A_{j})$
for all distinct $i$ and $j$.
Proof.
We proceed by induction on the number $k$ of homotopy classes. The statement is already seen to be true for $k\leq 2$. Assume now that $k>2$ and that the statement holds for all smaller values of $k$.
By the inductive hypothesis, there exist representatives $a_{1},\dots,a_{k-1}$ of $A_{1},\dots,A_{k-1}$, respectively, such that
$\lvert\pi_{0}(a_{i}\cap a_{j})\rvert=i(A_{i},A_{j})$
for all distinct $i,j\in\{1,\dots,k-1\}$. Choose a representative $a_{k}\in A_{k}$. It suffices to show that if
$\sum_{i=1}^{k-1}\lvert\pi_{0}(a_{i}\cap a_{k})\rvert>\sum_{i=1}^{k-1}i(A_{i},A_{k}),$
then there exists another representative $a_{k}^{\prime}\in A_{k}$ satisfying
$\sum_{i=1}^{k-1}\lvert\pi_{0}(a_{i}\cap a_{k}^{\prime})\rvert<\sum_{i=1}^{k-1}\lvert\pi_{0}(a_{i}\cap a_{k})\rvert$.
This will produce a representative of $A_{k}$ in minimal position with respect to all $A_{i}$ for $i<k$.
Case 1. Some $a_{i}$ is an essential simple closed curve.
Without loss of generality, assume $a_{k}$ is such a curve. Let $\mathcal{A}$ denote the set of all subarcs $b\subset a_{k}$ for which $b\cup c$ bounds a bigon with some subarc $c\subset a_{i}$. The set $\mathcal{A}$ is partially ordered by inclusion. Choose a minimal element $b_{0}\in\mathcal{A}$ and let $c_{0}\subset a_{i}$ be such that $b_{0}\cup c_{0}$ bounds a disk $D_{0}$. For each $j\neq i$, minimality of $b_{0}$ implies that every component of $a_{j}\cap D_{0}$ is an arc joining $b_{0}$ and $c_{0}$. (If a component is a circle or a point, one may shrink $D_{0}$ slightly to remove it.) Let $a_{k}^{\prime}$ be the simple closed curve obtained by homotoping $b_{0}$ to $c_{0}$. Then $\lvert a_{j}\cap a_{k}^{\prime}\rvert\leq\lvert a_{j}\cap a_{k}\rvert$ for all $j\neq i$, as required.
Case 2. None of $a_{1},\dots,a_{k}$ is an essential simple closed curve.
The argument is analogous. Suppose $a_{k}$ and $a_{i}$ form a bihedron for some $i$; that is, there exist disks $d_{i}\subset a_{i}$ and $d_{k}\subset a_{k}$ such that $d_{i}\cup d_{k}$ is a sphere bounding a 33–ball. Assume $d_{k}$ is minimal in the sense that no proper subset of $d_{k}$ forms a bigon or a bihedron with any $a_{j}$. Let $B_{0}$ denote the 33–ball bounded by $d_{i}\cup d_{k}$. By the inductive hypothesis, and by minimality of $d_{k}$, each intersection $a_{j}\cap B_{0}$ is a disjoint union of disks and cylinders connecting $d_{i}$ and $d_{k}$. Let $a_{k}^{\prime}$ be obtained from $a_{k}$ by homotoping $d_{k}$ to $d_{i}$. Then $\lvert\pi_{0}(a_{j}\cap a_{k}^{\prime})\rvert\leq\lvert\pi_{0}(a_{j}\cap a_{k})\rvert$ for all $j\neq i$. The case in which $a_{k}$ and $a_{i}$ form a bigon is treated similarly and will be omitted.
By iterating this procedure, we obtain a representative $a_{k}^{\prime\prime}\in A_{k}$ satisfying $\lvert\pi_{0}(a_{i}\cap a_{k}^{\prime\prime})\rvert=i(A_{i},A_{k})$ for all $i\neq k$. Hence, all pairs among $a_{1},\dots,a_{k-1},a_{k}^{\prime\prime}$ are in minimal position, completing the proof. ∎
3.2. Intersection number and Dehn twist
In this section, we analyze when two Dehn twists commute. If their representatives have disjoint supports, then the corresponding Dehn twists clearly commute. We are interested in the converse question: if two Dehn twists commute, must they admit representatives with disjoint supports?
Lemma 3.6.
Let $S$ be a homotopy class of either an essential sphere or an essential torus, and let $T$ be a homotopy class of an essential torus. If $i(S,\delta_{\vec{T}}^{k}(S))=0$ for all $k>0$, then we have $i(S,T)=0$.
Proof.
We investigate how essential spheres and essential tori evolve under iteration of a Dehn twist by examining the intersection between an essential surface and its image under a power of the twist. Let $T$ be a homotopy class of an essential torus with core $\gamma$. Let $k$ be a positive integer. Fix a homotopy class $S$ of an essential sphere with $i(S,T)>0$, and choose representatives $s\in S$ and $t\in T$ in minimal position.
For each component $c$ of $s\cap t$, exactly one of the following holds:
  1. (1)
    $c$ bounds a disk in $t$; or
  2. (2)
    $c$ represents a meridian of $t$.
The number of intersection circles of type (2) is equal to $i(\gamma,S)$. Consequently, the number of intersection circles of type (1) is $i(S,T)-i(\gamma,S)$. The local modification of a neighborhood of $c$ under the Dehn twist $\delta_{\vec{T}}^{k}$ depends essentially on whether $c$ is of type (1) or (2).
Figure 1. $c$ bounds a disk in $t$.
If (1) holds, that is, $c$ bounds a disk in $t$, then a neighborhood of $c$ in $s$ is stretched along the core $\gamma$ of $t$ under a positive power of the Dehn twist. As a result, this neighborhood intersects $s$ in
$k\,i(\gamma,S)+2$
circles. See Figure 1.
Figure 2. $c$ represents a meridian of $t$.
If (2) holds, that is, $c$ represents a meridian of $t$, then a neighborhood of $c$ in $s$ is stretched along the entire torus $t$ by a positive of the Dehn twist. In this case, the neighborhood intersects $s$ in
ki(S,T)1k\,i(S,T)-1
circles. See Figure 2.
Figure 3. The surfaces $s$ and $t$ intersect in three circles. In this configuration, there exists an essential torus homotopic to $\delta_{\vec{t}}(s)$ intersects $s$ in eight circles.
From the above local pictures, one obtains a corresponding global configuration. Figure 3 illustrates how these local modifications assemble into a global intersection pattern. Combining the contributions from each type of intersection circle, we obtain the following estimate: for every $k>0$,
$i(S,\delta_{\vec{T}}^{k}(S))$ $\geq(i(S,T)-i(\gamma,S))\,(k\,i(\gamma,S)+2)+i(\gamma,S)(k\,i(S,T)-1)$
$=k\,i(\gamma,S)\,(2\,i(S,T)-i(\gamma,S))+2\,i(S,T)-3\,i(\gamma,S).$
To justify this inequality, it suffices to verify that the resulting global configuration contains neither bigons nor bihedra. First, none of the local modifications described above produces a bihedron; consequently, no bihedron appears in the global picture. Next, suppose for contradiction that there exists a bigon between $s$ and $s_{k}$ where $s_{k}$ is a representative of $\delta_{\vec{T}}^{k}(S)$ locally described as the above. Then the arc of $s_{k}$ forming one side of the bigon can be decomposed into a concatenation of arcs, each of which lies entirely in either $s$ or $t$. It follows that there exists a bigon formed by $s$ and $t$. This contradicts the assumption that $s$ and $t$ are in minimal position. Therefore, no bigons occur between $s$ and $s_{k}$, and the inequality follows.
Thus $k\geq 2$ is a sufficient condition to guarantee that $i(S,\delta_{\vec{T}}^{k}(S))>0$. Equivalently, if $i(S,T)>0$ and $k\geq 2$, then
$i(S,\delta_{\vec{T}}^{k}(S))>0.$
This is the desired conclusion.
Suppose two homotopy classes $T$ and $T^{\prime}$ of essential tori have a positive intersection number, that is, $i(T,T^{\prime})>0$. Choose representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ in minimal position. We say an essential simple closed curve of $t$ is a meridian if it bounds a disk in $\mathbb{M}$. Then each intersection circle $c$ of $t$ and $t^{\prime}$ satisfies one of the following:
  1. (1)
    $c$ bounds a disk in each of $t$ and $t^{\prime}$;
  2. (2)
    $c$ bounds a disk in $t$ and represents a meridian of $t^{\prime}$;
  3. (3)
    $c$ represents a meridian of $t$ and bounds a disk in $t^{\prime}$; or
  4. (4)
    $c$ represents a meridian of both $t$ and $t^{\prime}$.
The type of each intersection circle in the above list is invariant under homotopy. More precisely, for any other choice of representatives of $T$ and $T^{\prime}$ in minimal position, each of the above types of intersection circles occurring in $t\cap t^{\prime}$ also occurs for the new representatives.
Let us examine how the local picture near an intersection circle changes after applying a positive power of the Dehn twist $\delta_{\vec{T}}^{k}$ to $T^{\prime}$. Fix representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ in minimal position, and let $c$ be a component of $t\cap t^{\prime}$.
Type (1). Suppose that $c$ bounds a disk in each of $t$ and $t^{\prime}$. Then there exist disks $D\subset t$ and $D^{\prime}\subset t^{\prime}$ with $\partial D=\partial D^{\prime}=c$ such that $D\cup D^{\prime}$ is an essential sphere; see the left-hand side of Figure 4. A representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ can be chosen so that, in a neighborhood of $c$, it has the local form shown on the right-hand side of Figure 4. In particular, this local modification creates at least two transverse intersection curves with $t^{\prime}$.
Figure 4. An intersection circle bounds a disk in each of $t$ and $t^{\prime}$.
Type (2). Suppose that $c$ bounds a disk in $t$ and represents a meridian of $t^{\prime}$. If $c$ appears as on the left-hand side of Figure 5, then a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side of Figure 5. In particular, in this neighborhood the surfaces $\delta_{\vec{T}}^{k}(t^{\prime})$ and $t^{\prime}$ intersect in at least two circles.
Figure 5. An intersection circle bounds a disk in $t$ and represents a meridian of $t^{\prime}$.
Type (3). Suppose that $c$ represents a meridian of $t$ and bounds a disk in $t^{\prime}$. Then $t$ and $t^{\prime}$ meet locally as on the left-hand side of Figure 6, and a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side. In this case the number of intersection circles of $t^{\prime}$ and $\delta_{\vec{T}}^{k}(t^{\prime})$ contained in a sufficiently small neighborhood of $c$ grows linearly with $k$. More precisely, this number is bounded below by $k\,i(T,T^{\prime})-1$.
Figure 6. An intersection circle represents a meridian of $t$ and bounds a disk in $t^{\prime}$.
Type (4). Suppose that $c$ represents a meridian of both $t$ and $t^{\prime}$. Then $t$ and $t^{\prime}$ meet locally as on the left-hand side of Figure 7, and a representative of $\delta_{\vec{T}}^{k}(T^{\prime})$ has the local form shown on the right-hand side. In this case the number of intersection circles of $t^{\prime}$ and $\delta_{\vec{T}}^{k}(t^{\prime})$ contained in a sufficiently small neighborhood of $c$ equals $k\,i(T,T^{\prime})$.
Figure 7. An intersection circle represents a meridian of both $t$ and $t^{\prime}$.
We now estimate $i(\delta_{\vec{T}}^{k}(T^{\prime}),T^{\prime})$. For each $k>0$, construct a representative $t_{k}^{\prime\prime}$ of $\delta_{\vec{T}}^{k}(T^{\prime})$ by applying the local modifications described in (1)–(4) in a neighborhood of every intersection circle of $t\cap t^{\prime}$. In general, the pair $(t_{k}^{\prime\prime},t^{\prime})$ need not be in minimal position: the local construction may create bigons between $t_{k}^{\prime\prime}$ and $t^{\prime}$. However, the number of such bigons is uniformly bounded (independent of $k$), since bigons can only occur in finitely many prescribed neighborhoods determined by the original configuration of $t\cap t^{\prime}$. Therefore, after removing all bigons by homotopy (i.e. by repeatedly applying the bigon criterion), the total number of intersection circles decreases by at most a uniform constant.
On the other hand, the local models in Figures 6 and 7 produce at least $k\,i(T,T^{\prime})^{2}-i(T,T^{\prime})$ intersection circles between $t_{k}^{\prime\prime}$ and $t^{\prime}$ whenever there is at least one intersection circle of type (3) or (4). Consequently, for $k$ sufficiently large, the intersection number remains positive after bigon removal, that is,
$i(\delta_{\vec{T}}^{k}(T^{\prime}),T^{\prime})>0$
for all sufficiently large $k$. This is the desired statement. ∎
One of the implications from Lemma 3.6 is Proposition 2.7.
Proof of Proposition 2.7.
Lemma 2.6 implies that a homotopy class of an (oriented) essential torus determines an isotopy class of a Dehn twist. Thus, to prove Proposition 2.7, it suffices to show that if two Dehn twist homeomorphisms are isotopic, then their corresponding essential tori are homotopic.
Suppose that $t$ and $t^{\prime}$ are essential tori that are not homotopic, however, in minimal position. We claim that there exists an essential sphere $s$ such that either
$i([s],[t])=0<i([s],[t^{\prime}])\quad\text{or}\quad i([s],[t])>0=i([s],[t^{\prime}]).$
Let $c$ be a meridian of $t$. If $D$ is an embedded disk with $t\cap D=c=\partial D$, then one obtains an essential sphere $s_{D}$ as a boundary component of a regular neighborhood of $t\cup D$. If there exists such a disk $D$ with $i([s_{D}],[t^{\prime}])>0$, then $s_{D}$ satisfies the claim.
Otherwise, that is, if $i([s_{D}],[t^{\prime}])=0$ for all $D$, then there exists an essential sphere $s$ that intersects $t$ exactly at $c$. By the bigon–bihedron criterion, $s$ and $t$ are in minimal position. Since $t^{\prime}$ is disjoint from $s$, this sphere $s$ satisfies the claim. In conclusion, the claim holds.
By Lemma 3.6, there exists $k>0$ such that either
$i([s],[\delta_{\vec{t}}^{k}(s)])=0<i([s],[\delta_{\vec{t}^{\prime}}^{k}(s)])\quad\text{or}\quad i([s],[\delta_{\vec{t}}^{k}(s)])>0=i([s],[\delta_{\vec{t}^{\prime}}^{k}(s)]).$
Therefore, $\delta_{\vec{t}}$ is not isotopic to $\delta_{\vec{t}^{\prime}}$. ∎
If two homeomorphisms have disjoint supports, then they commute. Consequently, their mapping classes also commute. In other words, mapping classes that admit representatives with disjoint supports must commute. However, in general, commutativity does not imply the existence of representatives with disjoint supports. On surfaces, it is known that Dehn twists commute if and only if they admit representatives with disjoint supports (see [5, Fact 3.6]). The following theorem shows that the same phenomenon holds for doubled handlebodies.
Theorem 3.7.
Let $T$ and $T^{\prime}$ be homotopy classes of essential tori. Then, we have $\delta_{\vec{T}},\delta_{\vec{T}^{\prime}}]=1$ for any orientations of $T$ and $T^{\prime}$ if and only if $i(T,T^{\prime})=0$.
Proof.
Suppose first that $i(T,T^{\prime})=0$. Then there exist representatives $t\in T$ and $t^{\prime}\in T^{\prime}$ which are disjoint. Since the supports of $\delta_{\vec{t}}$ and $\delta_{\vec{t}^{\prime}}$ are disjoint, these homeomorphisms commute. It follows that $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute.
Conversely, assume that $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute. Then for each $k>0$, we have
$\delta_{\delta_{\vec{T}^{\prime}}^{k}(\vec{T})}=\delta_{\vec{T}^{\prime}}^{k}\delta_{\vec{T}}\delta_{\vec{T}^{\prime}}^{-k}=\delta_{\vec{T}}.$
So one gets $\delta_{\vec{T}^{\prime}}^{k}(T)=T$ for every $k>0$. By Lemma 3.6, this implies that $i(T,T^{\prime})=0$. Therefore, $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute if and only if $i(T,T^{\prime})=0$. ∎
4. On the universal cover
Let $p:\widetilde{\mathbb{M}}\to\mathbb{M}$ be a universal cover. Any complete geodesic Riemannian metric on $\mathbb{M}$ induces Gromov hyperbolicity on $\widetilde{\mathbb{M}}$. So we obtain a Gromov boundary of $\widetilde{\mathbb{M}}$, denoted by $\partial\widetilde{\mathbb{M}}$. Two bi-infinite lines on $\widetilde{\mathbb{M}}$ are said to be equivalent if the Gromov–Hausdorff distance between them is finite. For a bi-infinite line $\tilde{c}$ and an essential sphere $\tilde{s}$ in $\widetilde{\mathbb{M}}$, we say $\tilde{c}$ penetrates $\tilde{s}$ if $\tilde{c}$ and $\tilde{s}$ intersect without bigon. For two integers $a$ and $b$, let $\operatorname{lcm}(a,b)$ denote the least common multiple of $a$ and $b$.
Lemma 4.1.
Let $c$ and $c^{\prime}$ be essential simple closed curves, and let $s$ be an essential sphere, which are in minimal position. Let $\tilde{c}$ and $\tilde{c}^{\prime}$ be lifts of $c$ and $c^{\prime}$, respectively. If $\tilde{c}$ is not equivalent to $\tilde{c}^{\prime}$, then the number of lifts of $s$ penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$ is at most $\operatorname{lcm}(i([c],[s]),i([c^{\prime}],[s]))$.
Proof.
If $\operatorname{lcm}(i([c],[s]),i([c^{\prime}],[s]))=0$, then $s$ is disjoint from either $c$ or $c^{\prime}$, that is, there is no lifts of $s$ penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$. So the statement holds in this case. Let us assume that $\operatorname{lcm}(i([c],[s]),i([c^{\prime}],[s]))>0$.
Suppose, for contradiction, that there are at least $\operatorname{lcm}(i([c],[s]),i([c^{\prime}],[s]))+1$ lifts of $s$, each of which penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$. Then there are two distinct lifts $\tilde{s}_{1}$ and $\tilde{s}_{2}$ of $s$ such that
$p(\tilde{c}\cap\tilde{s}_{1})=p(\tilde{c}\cap\tilde{s}_{2})$ and $p(\tilde{c}^{\prime}\cap\tilde{s}_{1})=p(\tilde{c}^{\prime}\cap\tilde{s}_{2})$.
If $w\in\mathbb{F}$ is an element that maps $\tilde{s}_{1}$ to $\tilde{s}_{2}$, then $w$ preserves $\tilde{c}$ and $\tilde{c}^{\prime}$. So $\tilde{c}$ and $\tilde{c}^{\prime}$ are equivalent, which contradicts the hypothesis. Therefore, the statement holds. ∎
A bi-infinite cylinder in $\widetilde{\mathbb{M}}$ is said to be essential if it does not bound a solid cylinder. Note that every lift of an essential torus is an essential cylinder. We also say a bi-infinite line $\tilde{c}$ penetrates an essential cylinder $\tilde{t}$ in $\widetilde{\mathbb{M}}$ if they intersect without bigon.
Proposition 4.2.
Let $c$ and $c^{\prime}$ be essential simple closed curves, and let $t$ be an essential torus, which are in minimal position. Let $\tilde{c}$ be a lift of $c$, and let $\tilde{c}^{\prime}$ be a lift of $c^{\prime}$. If $\tilde{c}$ is not equivalent to $\tilde{c}^{\prime}$, then the number of lifts of $t$ penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$ is at most $2\operatorname{lcm}(i([c],[t]),i([c^{\prime}],[t]))$.
Meanwhile, $\mathbb{F}$ acts properly discontinuously on $\widetilde{\mathbb{M}}$. We need the following property of $\mathbb{F}$ for a proof of Proposition 4.2.
Theorem 4.3 (Lyndon and Schützenberger [16]).
Let $x,y,z\in\mathbb{F}$ be nontrivial primitive elements. Suppose the equation $x^{a}y^{b}z^{c}=1$ holds for some $a,b,c\geq 2$. Then there exists an element $w\in\mathbb{F}$ such that $x,y,z\in\{w,w^{-1}\}$.
We say an element $g\in\mathbb{F}$ minimally preserves a bi-infinite line $\tilde{c}$ if $h$ is any element of $\mathbb{F}$ that preserves $\tilde{c}$, then one gets $h=g^{n}$ for some $n\in\mathbb{Z}$.
Proof of Proposition 4.2.
Suppose there are at least $2\operatorname{lcm}(i([c],[t]),i([c^{\prime}],[t]))+1$ lifts of $t$ which are penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$. Then there are three lifts $\tilde{t}_{1}$, $\tilde{t}_{2}$ and $\tilde{t}_{3}$ of $t$ such that
$p(\tilde{c}\cap\tilde{t}_{1})=p(\tilde{c}\cap\tilde{t}_{2})=p(\tilde{c}\cap\tilde{t}_{3})$ and $p(\tilde{c}^{\prime}\cap\tilde{t}_{1})=p(\tilde{c}^{\prime}\cap\tilde{t}_{2})=p(\tilde{c}^{\prime}\cap\tilde{t}_{3})$.
Let $g$ and $g^{\prime}$ be elements of $\mathbb{F}$ minimally preserving $\tilde{c}$ and $\tilde{c}^{\prime}$, respectively. Let us take signs for $g$ and $g^{\prime}$ to satisfy that
$g^{m+n}(\tilde{t}_{1})=g^{n}(\tilde{t}_{2})=\tilde{t}_{3}$ and $(g^{\prime})^{m^{\prime}+n^{\prime}}(\tilde{t}_{1})=(g^{\prime})^{n^{\prime}}(\tilde{t}_{2})=\tilde{t}_{3}$
for some positive integers $m,n,m^{\prime},n^{\prime}$. Then $(g^{\prime})^{m^{\prime}}g^{-m}$ and $(g^{\prime})^{m^{\prime}+n^{\prime}}g^{-(m+n)}$ preserve $\tilde{t}_{2}$ and $\tilde{t}_{3}$, respectively. If $(g^{\prime})^{m^{\prime}}g^{-m}$ is trivial, then $\tilde{c}$ and $\tilde{c}^{\prime}$ are equivalent, so this contradicts the hypothesis. Otherwise, there exists a nontrivial element $h_{2}\in\mathbb{F}$ that preserves $\tilde{t}_{2}$ such that $h_{2}=(g^{\prime})^{m^{\prime}}g^{-m}$.
Similarly, since $g^{-(m+n)}(g^{\prime})^{m^{\prime}+n^{\prime}}$ is nontrivial, there exists a nontrivial element $h_{3}\in\mathbb{F}$ that preserves $\tilde{t}_{3}$ such that $h_{3}=(g^{\prime})^{m^{\prime}+n^{\prime}}g^{-(m+n)}$. If $h_{3}$ is conjugate to $h_{2}$, then this implies that $\tilde{c}$ and $\tilde{c}^{\prime}$ are equivalent. Otherwise, $h_{3}$ can be expressed as a nontrivial power of some primitive element $h$. That is, we get an equation $h^{n}=(g^{\prime})^{m^{\prime}+n^{\prime}}g^{-(m+n)}$ for some $n\geq 2$. By Theorem 4.3, $h,g,g^{\prime}$ are powers of a common element. That is, $\tilde{c}$ and $\tilde{c}^{\prime}$ are equivalent. This is a contradiction. Therefore, there are at most $2\operatorname{lcm}(i([c],[t]),i([c^{\prime}],[t]))$ lifts of $t$ that are penetrated by both $\tilde{c}$ and $\tilde{c}^{\prime}$. ∎
Proposition 4.4.
Let $c$ be an essential simple closed curve, and let $t$ and $t^{\prime}$ be essential tori in minimal position. Let $\tilde{t}_{-1},\tilde{t}_{0},\tilde{t}_{1}$ be a consecutive sequence of lifts of $t$ penetrated by some lift $\tilde{c}$ of $c$. For each $i$, let $\tilde{c}_{i}$ be the core of $\tilde{t}_{i}$. Then the number of lifts of $t^{\prime}$ penetrated by $\tilde{c}_{0}$ that separate $\tilde{c}_{-1}$ from $\tilde{c}_{1}$ is at most $i([c],[t^{\prime}])$.
Proof.
[…]
5. Compatible collection and fellow traveler
In this section, we introduce compatible collections and the notion of fellow travelers.
5.1. Compatible collection
A homotopy class of an essential simple closed curve is said to penetrate a homotopy class of an essential torus if their intersection number is positive. Given two homotopy classes $T$ and $T^{\prime}$ of essential tori, if the core of $T$ penetrates $T^{\prime}$, we simply say that $T$ penetrates $T^{\prime}$. We say that $T$ and $T^{\prime}$ co-penetrate if they penetrate each other. A finite collection $\{T_{1},\dots,T_{k}\}$ of homotopy classes of essential tori is called compatible if their cores are pairwise distinct and every pair that intersects also co-penetrates.
5.2. Fellow travelers
Let $V=\{T_{1},\dots,T_{k}\}$ be a compatible collection of homotopy classes of essential tori, and let $\Gamma$ be the coincidence graph of $V$. Thus the vertices of $\Gamma$ are $T_{1},\dots,T_{k}$, and $T_{i}$ is joined to $T_{j}$ by an edge if and only if $i(T_{i},T_{j})=0$. We adopt standard graph-theoretic notation: for instance, $\operatorname{lk}(T_{i})$ denotes the link of $T_{i}$ and $\operatorname{st}(T_{i})$ its star.
For a homotopy class $\gamma$ of an essential simple closed curve, a lift $\tilde{\gamma}$ of $\gamma$ means an equivalence class of lifts of representatives of $\gamma$. Likewise, for a homotopy class $T$ of an essential torus, a lift $\tilde{T}$ of $T$ refers to an equivalence class of lifts of representatives of $T$. We say that $\tilde{\gamma}$ penetrates $\tilde{T}$ if any representative of $\tilde{\gamma}$ intersects any representative of $\tilde{T}$.
For a positive integer $N$, we say a homotopy $\gamma$ of an essential simple closed curve NN–fellow-travels with $T_{i}$ (with respect to $\Gamma$) if $\gamma$ penetrates $T_{i}$ and there are lifts $\tilde{\gamma}$ of $\gamma$ and $\tilde{\gamma}_{i}$ of $\gamma_{i}$ such that for each $T_{j}\in\Gamma\setminus\operatorname{st}(T_{i})$, at least $N$ lifts of $T_{j}$ are penetrated by both $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$. For each $i$, let $\mathcal{FT}\!_{N}(T_{i},\Gamma)$ denote the collection of NN–fellow travelers of $T_{i}$ with respect to $\Gamma$. By definition, the sequence
$\mathcal{FT}\!_{1}(T_{i},\Gamma)\supseteq\mathcal{FT}\!_{2}(T_{i},\Gamma)\supseteq\mathcal{FT}\!_{3}(T_{i},\Gamma)\supseteq\dots$
is satisfied. Define
$\Delta(\Gamma):=\max_{i,j,\ell=1,\dots,k}\operatorname{lcm}(i(\gamma_{i},T_{j}),\,i(\gamma_{\ell},T_{j})),$
where $\operatorname{lcm}$ denotes the least common multiple. Then the following holds.
Lemma 5.1.
Let $\gamma$ be a homotopy class of an essential simple closed curve that $(4\,\Delta(\Gamma)+3)$–fellow-travel with $T_{i}$. Then there exist lifts $\tilde{\gamma}$ of $\gamma$ and $\tilde{\gamma}_{i}$ of $\gamma_{i}$ such that the following condition is satisfied: for every $T_{j}\in\Gamma\setminus\operatorname{st}(T_{i})$, there exist pairwise distinct lifts $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}$ of $T_{j}$ with the property that
  1. (1)
    each of $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}$ is penetrated by both $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$, and
  2. (2)
    whenever $\tilde{T}_{\ell}$ is a lift of some $T_{\ell}\in\operatorname{lk}(T_{i})$ penetrated by $\tilde{\gamma}$, then the core of $\tilde{T}_{\ell}$ does not penetrate any of $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}$.
Proof.
By the hypothesis, there exist lifts $\tilde{T}_{j}^{(1)},\dots,\tilde{T}_{j}^{(J)}$ of $T_{j}$, with $J\geq 4\,\Delta(\Gamma)+3$, each of which is penetrated by both $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$. Fix an orientation on $\tilde{\gamma}$; we may assume that the lifts $\tilde{T}_{j}^{(1)},\dots,\tilde{T}_{j}^{(J)}$ are indexed in the order in which $\tilde{\gamma}$ meets them.
Now choose a lift $\tilde{T}_{\ell}$ of some $T_{\ell}\in\operatorname{lk}(T_{i})$ that is penetrated by $\tilde{\gamma}$, and let $\tilde{\gamma}_{\ell}$ denote its core. By Proposition 4.2, at most $2\,\operatorname{lcm}(i(\gamma_{i},T_{j}),i(\gamma_{\ell},T_{j}))$ lifts of $T_{j}$ can be penetrated by both $\tilde{\gamma}_{\ell}$ and $\tilde{\gamma}_{i}$. Since $\Delta(\Gamma)$ bounds all such least common multiples, it follows that $\tilde{T}_{j}^{(2\,\Delta(\Gamma)+1)},\dots,\tilde{T}_{j}^{(J-2\,\Delta(\Gamma))}$ are not penetrated by $\tilde{\gamma}_{\ell}$.
Choose three pairwise distinct lifts $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}\in\{\tilde{T}_{j}^{(2\,\Delta(\Gamma)+1)},\dots,\tilde{T}_{j}^{(J-2\,\Delta(\Gamma))}\}$. By the first paragraph, each of them is penetrated by both $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$. Furthermore, the second paragraph implies that none of $\tilde{T}_{j},\tilde{T}_{j}^{\prime}$, or $\tilde{T}_{j}^{\prime\prime}$ is penetrated by the core of any lift of any $T_{\ell}\in\operatorname{lk}(T_{i})$. This establishes the desired conclusion. ∎
5.3. Dehn twists and fellow travelers
For each $i$, we write $\delta_{i}$ for the Dehn twist $\delta_{\vec{T}_{i}}$ along $T_{i}$, choosing some orientation of $T_{i}$. Define
$M(\Gamma):=\max_{i,j=1,\dots,k}i(\gamma_{i},T_{j}).$
The next proposition shows that, after applying a suitable power of $\delta_{i}$ to certain homotopy classes of essential simple closed curves, the resulting curve fellow-travels with $T_{i}$ for a long time. As a consequence, each set $\mathcal{FT}\!_{4\,\Delta(\Gamma)+3}(T_{i},\Gamma)$ contains infinitely many homotopy classes of essential simple closed curves.
Proposition 5.2.
If an integer $n$ satisfies $\lvert n\rvert\geq 5\,M(\Gamma)+8$, then we have
$\delta_{j}^{n}(\mathcal{FT}\!_{3}(T_{i},\Gamma)\cup\{\gamma_{i}\})\subseteq\mathcal{FT}\!_{4\,\Delta(\Gamma)+3}(T_{j},\Gamma)$
for all co-penetrating pairs $T_{i}$ and $T_{j}$.
Proof.
Let $\gamma\in\mathcal{FT}\!_{3}(T_{i},\Gamma)\cup\{\gamma_{i}\}$ be given. By definition, there are lifts $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$ of $\gamma$ and $\gamma_{i}$, respectively, such that they penetrate three consecutive lifts $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}$ of $T_{j}$. Choose $T_{\ell}$ such that $T_{j}$ and $T_{\ell}$ co-penetrate. Let $\tilde{\gamma}_{j},\tilde{\gamma}_{j}^{\prime},\tilde{\gamma}_{j}^{\prime\prime}$ be the cores of $\tilde{T}_{j},\tilde{T}_{j}^{\prime},\tilde{T}_{j}^{\prime\prime}$, respectively. Let $\tilde{\delta}_{j}$ be a lift of $\delta_{j}$ that preserves $\tilde{T}_{j}$ and $\tilde{T}_{j}^{\prime}$. Then there exists a nontrivial element $g\in\mathbb{F}$ such that $\tilde{\delta}_{j}^{n}(\tilde{\gamma}_{j}^{\prime\prime})=g^{n}\tilde{\gamma}_{j}^{\prime\prime}$ where $g$ preserves $\tilde{\gamma}_{j}^{\prime}$.
We claim that the number of lifts of $T_{\ell}$ that are penetrated by $\tilde{\gamma}_{j}^{\prime}$ and that separate $\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$ is at least
$(\lvert n\rvert-3)\,i(\gamma_{j},T_{\ell}).$
Without loss of generality, assume $n>0$; the case $n<0$ follows by the same argument applied to $g^{-1}$ in place of $g$. Let $\tilde{T}_{\ell}$ be any lift of $T_{\ell}$ penetrated by $\tilde{\gamma}_{j}^{\prime}$.
If $g^{a}\tilde{T}_{\ell}$ is never penetrated by $\tilde{\gamma}_{j}^{\prime\prime}$ for any $a\in\mathbb{Z}$, then there exists $a_{0}\in\mathbb{Z}$ such that $g^{a_{0}}\tilde{T}_{\ell}$ separates $g^{-b}\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{c}\tilde{\gamma}_{j}^{\prime\prime}$ for all $b,c>0$. Consequently,
$g^{a_{0}}\tilde{T}_{\ell},\ g^{a_{0}+1}\tilde{T}_{\ell},\ \dots,\ g^{a_{0}+n-1}\tilde{T}_{\ell}$
all separate $\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$. If instead $g^{a_{1}}\tilde{T}_{\ell}$ is penetrated by $\tilde{\gamma}_{j}^{\prime\prime}$ for some $a_{1}\in\mathbb{Z}$, then by Proposition 4.2 the lift $g^{a_{1}}\tilde{T}_{\ell}$ separates $g^{-b-1}\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{c+1}\tilde{\gamma}_{j}^{\prime\prime}$ for all $b,c>0$. It follows that
$g^{a_{1}+2}\tilde{T}_{\ell},\ g^{a_{1}+3}\tilde{T}_{\ell},\ \dots,\ g^{a_{1}+n-2}\tilde{T}_{\ell}$
all separate $\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$.
Since the lifts of $T_{\ell}$ penetrated by $\tilde{\gamma}_{j}^{\prime}$ consist of exactly $i(\gamma_{j},T_{\ell})$ distinct $\langle g\rangle$–orbits, the above argument shows that the total number of such separating lifts is at least
$(n-3)\,i(\gamma_{j},T_{\ell})=(|n|-3)\,i(\gamma_{j},T_{\ell}).$
This proves the claim.
Let $\mathcal{T}$ be the set of lifts of $T_{\ell}$ that are penetrated by $\tilde{\gamma}_{j}^{\prime}$ and that separate $\tilde{\gamma}_{j}^{\prime\prime}$ from $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$. By Proposition 4.2, at most $2\,i(\gamma_{j},T_{\ell})$ lifts in $\mathcal{T}$ are penetrated by $\tilde{\gamma}_{j}$. If a lift in $\mathcal{T}$ is not penetrated by $\tilde{\gamma}_{j}$, then it separates $\tilde{\gamma}_{j}$ from either $\tilde{\gamma}_{j}^{\prime\prime}$ or $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$. By Proposition 4.4, the number of such lifts separating $\tilde{\gamma}_{j}$ from $\tilde{\gamma}_{j}^{\prime\prime}$ is at most $i(\gamma_{i},T_{\ell})$. Therefore, the number of lifts in $\mathcal{T}$ that separate $\tilde{\gamma}_{j}$ from $g^{\,n}\tilde{\gamma}_{j}^{\prime\prime}$ is at least
$(|n|-3)\,i(\gamma_{j},T_{\ell})\;-\;2\,i(\gamma_{j},T_{\ell})\;-\;i(\gamma_{i},T_{\ell})\;\geq\;4\,\Delta(\Gamma)+3.$
[…]
Proposition 5.3.
If $T_{i},T_{j_{1}},\dots,T_{j_{m}}$ are pairwise distinct and pairwise disjoint, i.e., $i(T_{a},T_{b})=0$ for all $a,b\in\{i,j_{1},\dots,j_{m}\}$, then we have
$\delta_{j_{m}}^{n_{m}}\dots\delta_{j_{1}}^{n_{1}}(\mathcal{FT}\!_{4\,\Delta(\Gamma)+3}(T_{i},\Gamma))\subseteq\mathcal{FT}\!_{3}(T_{i},\Gamma)$
for all $n_{1},\dots,n_{m}\in\mathbb{Z}$.
Proof.
Choose $\gamma\in\mathcal{FT}\!_{4\,\Delta(\Gamma)+3}(T_{i},\Gamma)$. By Lemma 5.1, for each $T_{\ell}\in\Gamma\setminus\operatorname{st}(T_{i})$, there are three lifts of $T_{\ell}$, denoted by $\tilde{T}_{\ell},\tilde{T}_{\ell}^{\prime},\tilde{T}_{\ell}^{\prime\prime}$, consecutively penetrated by $\tilde{\gamma}$ and $\tilde{\gamma}_{i}$ that are not penetrated by any core of a lift of $T_{j_{1}},\dots,T_{j_{\ell}}$.
For each $a\in\{1,\dots,\ell\}$, let $\tilde{\delta}_{j_{a}}$ be the lift of $\delta_{j_{a}}$ that fixes $\tilde{\gamma}_{i}$ pointwise. Let $\mathcal{O}$ be the collection of the lifts of $T_{j_{1}},\dots,T_{j_{\ell}}$ penetrated by $\tilde{\gamma}$. Note $\mathcal{O}$ can be ordered by an orientation of $\tilde{\gamma}$. Let $\tilde{T}^{\prime}$ and $\tilde{T}^{\prime\prime}$ be the consecutive elements of $\mathcal{O}$ such that the connected component of $\widetilde{\mathbb{M}}\setminus\bigsqcup_{\tilde{T}\in\mathcal{O}}\tilde{T}$ that contains $\tilde{\gamma}$ is bounded by $\tilde{T}^{\prime}\cup\tilde{T}^{\prime\prime}$.
[…]
Remark 5.4.
It is crucial that the cores of $T_{i},T_{j_{1}},\dots,T_{j_{\ell}}$ are pairwise distinct in order for Proposition 5.3 to hold. Otherwise, the composition $\delta_{j_{\ell}}^{n_{\ell}}\dots\delta_{j_{1}}^{n_{1}}$ may coincide with a power of $\delta_{i}$ which could send an element of $\mathcal{FT}\!_{4\,\Delta(\Gamma)+3}(T_{i},\Gamma)$ outside of $\mathcal{FT}\!_{3}(T_{i},\Gamma)$.
5.4. Right-angled Artin group generated by powers of Dehn twists
In this section, we look into a specific collection of essential tori whose sufficiently large powers of Dehn twists generate a right-angled Artin group. Note that there exist finitely many Dehn twists that generate $\operatorname{Out}(\mathbb{F})$, which is more complicate than right-angled Artin groups. This section is inspired by Koberda [14] and Seo [17].
Let us remind some definitions. A collection $V=\{T_{1},\dots,T_{k}\}$ of homotopy classes of essential tori is said to be compatible if their cores are pairwise distinct and whenever $i(T_{i},T_{j})>0$, they co-penetrate. The coincidence graph of $V$ is a simple graph whose vertices are $T_{1},\dots,T_{k}$ where $T_{i}$ and $T_{j}$ are joined by an edge if and only if $i(T_{i},T_{j})=0$.
Theorem 5.5.
Let $V=\{T_{1},\dots,T_{k}\}$ be a compatible collection of essential tori with cores $\gamma_{1},\dots,\gamma_{k}$, and let $\Gamma$ be the coincidence graph of $V$. For each $i$, let $\delta_{i}=\delta_{\vec{T}_{i}}$ be a Dehn twists along $T_{i}$. Then for all
$\lvert n_{1}\rvert,\dots,\lvert n_{k}\rvert\geq 5\,\max_{i\neq j}{i(\gamma_{i},T_{j})}+8,$
the homomorphism
$\Phi:A(\Gamma)\to\operatorname{Out}(\mathbb{F})$
defined by
$T_{i}\mapsto\delta_{i}^{n_{i}}$
is injective.
Proof.
We say $\Gamma$ is centerless if there is no complete subgraph $K$ such that $\Gamma\cong K*\Gamma^{\prime}$ for some $\Gamma^{\prime}\subset\Gamma$. First, we handle the centerless case of Theorem 5.5. Let us suppose that $\Gamma$ is centerless.
For each nontrivial element $x\in A(\Gamma)$, the support of $x$, denoted $\operatorname{supp}(x)$, is the subset of $V(\Gamma)$ such that $\operatorname{supp}(x)=\operatorname{supp}(w)$ for some reduced word $w$ representing $x$. This is well-defined: if both reduced words $w_{1}$ and $w_{2}$ represent the same element $x\in A(\Gamma)$, then $\operatorname{supp}(w_{1})=\operatorname{supp}(w_{2})$.
An element of $A(\Gamma)$ is said to be central if its support forms a complete subgraph of $\Gamma$. Every element of $A(\Gamma)$ can be expressed as a product of central elements, which is called a central form. A central form $x=x_{k}\cdots x_{1}$ is said to be in left greedy normal form if for each $i<k$ and each $v\in\operatorname{supp}(x_{i})$, there exists $u\in\operatorname{supp}(x_{i+1})$ such that $u,v]\neq 1$.
Since $\Gamma$ is centerless, there exists a vertex $T_{0}\in\Gamma\setminus\operatorname{st}(\operatorname{supp}(x_{1}))$. So there exists a vertex $T_{i_{1}}\in\operatorname{supp}(x_{1})$ such that $T_{i_{1}},T_{0}]\neq 1$. By Propositions 5.2 and 5.3, we have
$\Phi(x_{1})(\gamma_{0})\in\mathcal{FT}\!_{3}(T_{i_{1}},\Gamma).$
where $\gamma_{0}$ is the core of $T_{0}$. Inductively, one can show
$\Phi(x_{j}\dots x_{1})(\gamma_{0})$ $=\Phi(x_{j})(\Phi(x_{j-1}\dots x_{1})(\gamma_{0}))$
$\in\Phi(x_{j})(\mathcal{FT}\!_{3}(T_{i_{j-1}},\Gamma))\subseteq\mathcal{FT}\!_{3}(T_{i_{j}},\Gamma).$
for each $j=2,\dots,k$.
Since $\Phi(x)(\gamma_{0})\in\mathcal{FT}\!_{3}(T_{i_{k}},\Gamma)$ and $\gamma_{0}\notin\mathcal{FT}\!_{3}(T_{i_{k}},\Gamma)$, it follows that $\Phi(x)$ is nontrivial. Hence, the homomorphism $\Phi$ is injective.
Suppose that $\Gamma$ decomposes as a join $K*\Gamma^{\prime}$, where $K$ is a nonempty complete subgraph and $\Gamma^{\prime}$ is a (possibly empty) centerless subgraph. Then the center $\mathcal{Z}(A(\Gamma))$ is naturally identified with $\mathbb{Z}^{K}$, and the quotient $A(\Gamma)/\mathcal{Z}(A(\Gamma))$ is isomorphic to $A(\Gamma^{\prime})$. If $\ker(\Phi)$ is nontrivial, the argument from the previous subsection implies that
$\ker(\Phi)\ \leq\ \mathcal{Z}(A(\Gamma))\ \cong\ \mathbb{Z}^{K}.$
Choose an essential torus $T_{0}$ that co-penetrates every $T\in\Gamma$. Then the compatible graph of $V\cup\{T_{0}\}$ is centerless. Consequently, there exists $N^{\prime}>0$ such that for all integers $n_{0},n_{1},\dots,n_{k}\geq N^{\prime}$, the subgroup
$\langle\,\delta_{\vec{T}_{0}}^{n_{0}},\,\delta_{1}^{n_{1}},\,\dots,\,\delta_{k}^{n_{k}}\,\rangle$
is isomorphic to the right-angled Artin group
$A(\Gamma\sqcup\{T_{0}\})\cong A(\Gamma)*\mathbb{Z}\cong(A(\Gamma^{\prime})\times\mathbb{Z}^{K})*\mathbb{Z}.$
In particular, the quotient $\mathcal{Z}(A(\Gamma))/\ker(\Phi)$ contains a free-abelian subgroup of rank $\lvert K\rvert$, which forces $\ker(\Phi)$ to be trivial. Therefore, $\Phi$ is injective. ∎
6. The space of factorial curves
An essential simple closed curve is called factorial if it represents a $1$-free factor of $\mathbb{F}$. Let $\mathcal{C}$ denote the set of all homotopy classes of factorial curves. Let $\mathbb{R}_{+}$ denote the set of nonnegative real numbers, i.e., $0,\infty)$. Define a map
$i_{*}:\mathcal{C}\to\mathbb{R}_{+}^{\mathcal{S}},\qquad i_{*}(\gamma)(S)=i(\gamma,S)$
for all $\gamma\in\mathcal{C}$ and $S\in\mathcal{S}$. Let $P(\mathbb{R}_{+}^{\mathcal{S}})$ denote the projectivization of $\mathbb{R}_{+}^{\mathcal{S}}$, and let
$p:\mathbb{R}_{+}^{\mathcal{S}}\setminus\{0\}\to P(\mathbb{R}_{+}^{\mathcal{S}})$
be the natural projection. Since the image of $i_{*}$ contains no zero vector, the composition $p\circ i_{*}$ is well-defined.
Proposition 6.1.
The composition $p\circ i_{*}:\mathcal{C}\to P(\mathbb{R}_{+}^{\mathcal{S}})$ is injective.
Proof.
Assume that $\gamma,\gamma^{\prime}\in\mathcal{C}$ satisfy
$(p\circ i_{*})(\gamma)=(p\circ i_{*})(\gamma^{\prime}).$
Then there exists $\lambda>0$ such that
$i_{*}(\gamma)=\lambda\,i_{*}(\gamma^{\prime}).$
Since $\gamma$ corresponds to a free factor, there exists a separating sphere $s$ disjoint from a representative $c\in\gamma$ such that the component $C$ of $\mathbb{M}\setminus s$ containing $c$ is homeomorphic to the once–punctured doubled solid torus $(\mathbb{S}^{1}\times\mathbb{S}^{2})\setminus\{x\}$ for some point $x\in\mathbb{S}^{1}\times\mathbb{S}^{2}$. This region $C$ contains a unique isotopy class $S^{\prime}$ of essential spheres not isotopic to $s$. Choose a representative $s^{\prime}\in S^{\prime}$ such that $c$ intersects $s^{\prime}$ in exactly one point. Thus,
$i(\gamma,[s])=0,\qquad i(\gamma,S^{\prime})=\lvert c\cap s^{\prime}\rvert=1.$
By proportionality, it follows that
$i(\gamma^{\prime},[s])=0,\qquad i(\gamma^{\prime},S^{\prime})=\frac{1}{\lambda}>0.$
So $\gamma^{\prime}$ has a representative $c^{\prime}$ disjoint from $s$. Because $s$ is separating, the curve $c^{\prime}$ lies entirely inside $C$. Algebraically, the free factor determined by $c^{\prime}$ is contained in the free factor determined by $C$. Since $c^{\prime}$ is factorial, these free factors must be equal. Consequently, $c^{\prime}$ is homotopic to $c$, and hence $\gamma=\gamma^{\prime}$. ∎
By Proposition 6.1, we may regard $\mathcal{C}$ as a subspace of $P(\mathbb{R}_{+}^{\mathcal{S}})$.
Proposition 6.2.
The closure of $(p\circ i_{*})(\mathcal{C})$ is compact.
Before proving Proposition 6.2, we require the following technical lemma. This result is analogous to [6, Lemma 4.2].
Lemma 6.3.
Let $\Sigma$ be a maximal sphere system of $\mathbb{M}_{g}$. Then we have
$i(\gamma,S)\leq(2g-2)\,i(\gamma,\Sigma)\,i(S,\Sigma)$
for all $\gamma\in\mathcal{C}$ and $S\in\mathcal{S}$.
Proof.
Choose representatives $c\in\gamma$, $s\in S$, and $\sigma\in\Sigma$ that intersect pairwise minimally. Cutting $\mathbb{M}_{g}$ along $\sigma$ yields 2g22g-2 thrice-punctured spheres. For each such component $N$, the intersection $s\cap N$ is a disjoint union of annuli, pairs of pants, and disks, with at most $i(S,\Sigma)$ components. Similarly, $c\cap N$ is a disjoint union of at most $i(\gamma,\Sigma)$ arcs. Each of these arcs intersects $s$ in at most $i(S,\Sigma)$ points. So the number of intersection points of $c$ and $s$ inside $N$ is at most $i(\gamma,\Sigma)\,i(S,\Sigma)$. Summing over all thrice-punctured components gives the desired inequality. ∎
The following proof is analogous to [6, Proposition 4.1].
Proof of Proposition 6.2.
Fix an isotopy class of a maximal sphere system $\Sigma$. By Lemma 6.3, the set $\{i_{*}(\gamma)/i(\gamma,\Sigma)\mid\gamma\in\mathcal{C}\}$ is a subset of the bounded set
$\{f\in\mathbb{R}_{+}^{\mathcal{S}}\mid f(S)\leq(2g-2)\,i(S,\Sigma)\text{ for all }S\in\mathcal{S}\}.$
By Tychonoff’s theorem, the closure of $\{i_{*}(\gamma)/i(\gamma,\Sigma)\mid\gamma\in\mathcal{C}\}$ is compact. Consequently, the closure of $(p\circ i_{*})(\mathcal{C})$ is compact. ∎
The following is motivated by [6, Proposition A.1].
Lemma 6.4.
Let $T$ be a homotopy class of an essential torus with a core $\gamma$. If $\alpha$ is a homotopy class of a factorial curve and $S$ is an isotopy class of an essential sphere, then the following holds for all $n\in\mathbb{Z}$:
$\lvert i(\delta_{\vec{T}}^{n}(\alpha),S)-\lvert n\rvert\,i(\alpha,T)\,i(\gamma,S)\rvert\leq i(\alpha,S).$
Proof.
Let $a\in\alpha$, $t\in T$ and $s\in S$ be representatives in minimal position. Let $N(t)$ be a regular neighborhood of $t$ which is disjoint from $a\cap s$. Let $a_{n}$ be a representative of $\delta_{\vec{T}}^{n}(\alpha)$ that follows $a$ on the outside of $N(t)$ such that each component of $a_{n}\cap N(t)$ minimally intersect with $s$.
Then the number of components of $a_{n}\cap N(t)$ is exactly $i(\alpha,T)$. Furthermore, each component intersects with $s$ in $(\lvert n\rvert\,i(\gamma,S))$-times. So we obtain the following.
$i(\delta_{\vec{T}}^{n}(\alpha),S)$ $\leq\lvert a_{n}\cap s\rvert=\lvert(a_{n}-N(t))\cap s\rvert+\lvert(a_{n}\cap N(t))\cap s\rvert$
$=\lvert a\cap s\rvert+\lvert i(\alpha,T)\rvert\,\lvert n\rvert\,i(\gamma,S)$
$=i(\alpha,S)+\lvert n\rvert\,i(\alpha,T)\,i(\gamma,S)$
Suppose $a_{n}$ and $s$ form a bigon. Then there exists an arc $a^{\prime}\subset a_{n}$ and $b^{\prime}\subset s$ such that $a^{\prime}\cup b^{\prime}$ bounds a disk. Since $a$ and $s$ do not form a bigon, $a^{\prime}$ is not contained in $a$, that is, $a^{\prime}$ and $a_{n}\cap N(t)$ intersect. Any component of $a_{n}\cap N(t)$ cannot belong to $a^{\prime}$ because this minimally intersects with $s$. So we can decompose $a^{\prime}$ into two arcs $a^{\prime\prime}$ and $a^{\prime\prime\prime}$ such that $a^{\prime\prime}\in a$ and $a^{\prime\prime\prime}\in a_{n}\cap N(t)$.
By the above, reducing a bigon by homotoping $a_{n}$ induces the decrease of the number of $(a\cap a_{n})\cap s$. So there are at most $\lvert a\cap s\rvert$ reductions of bigons to make $a_{n}$ and $s$ become in minimal position. Since each bigon makes two intersection points, we get
$i(\delta_{\vec{T}}^{n}(\alpha),S)$ $\geq\lvert a_{n}\cap s\rvert-2\,\lvert a\cap s\rvert$
$=\lvert n\rvert\,i(\alpha,T)\,i(\gamma,S)-i(\alpha,S)$
Therefore, the inequality of the statement holds. ∎
The above lemma immediately induces the following proposition.
Theorem 6.5.
Let $T$ be a homotopy class of an essential torus with the core $\gamma$. Let $\alpha$ be a homotopy class of a factorial curve. Then, we have $i(\alpha,T)>0$ if and only if $(p\circ i_{*})(\delta_{\vec{T}}^{n}(\alpha))$ converges to $(p\circ i_{*})(\gamma)$ as $n\to\infty$.
Proof.
If $i(\alpha,T)=0$, then we have $\delta_{\vec{T}}^{n}(\alpha)=\alpha$ for all $n\in\mathbb{Z}$. Otherwise, we have
$\left\lvert\frac{1}{\lvert n\rvert\,i(\alpha,T)}\,i(\delta_{\vec{T}}^{n}(\alpha),S)-i(\gamma,S)\right\rvert\leq\frac{1}{\lvert n\rvert\,i(\alpha,T)}\,i(\alpha,S)$
for each $S\in\mathcal{S}$, so $(p\circ i_{*})(\delta_{\vec{T}}^{n}(\alpha))$ converges to $(p\circ i_{*})(\gamma)$ as $n\to\infty$. ∎
By Theorem 6.5, a core of a homotopy class of an essential torus is contained in the closure of $(p\circ i_{*})(\mathcal{C})$. Note a core of a homotopy class of an essential torus may not be a factorial curve. That is, $\overline{(p\circ i_{*})(\mathcal{C})}$ contains homotopy classes of essential simple closed curves that is not factorial.